What are the steps to answer:
4|2y-7|+5=9

Answer by fractalier(6550)

First we isolate the absolute value portion, so we do this...
4|2y-7| + 5 = 9
4|2y-7| = 4
|2y-7| = 1
Then we split it into two pieces and solve them separately
2y - 7 = 1 and 2y - 7 = -1
2y = 8 and 2y = 6
y = 4 and y = 3